OpenStudy (queenbb):

(Willl fan and medal) Which ordered pair is the solution to the system of equations? {x+y=−1 {3x−y=21 (10, −11) (7, −8) (5, −6) (5.5, −6.5)

OpenStudy (queenbb):

i need help

OpenStudy (daniyil_the_spy):

ok ill help

OpenStudy (queenbb):

yay

OpenStudy (mecharv):

Check your question... something seems off

OpenStudy (queenbb):

fixed it

OpenStudy (daniyil_the_spy):

x+y=−13x−y=21 x+y=−13x−y=21

OpenStudy (queenbb):

that what

OpenStudy (daniyil_the_spy):

both are not linear

OpenStudy (queenbb):

sorry i had to edit the question

OpenStudy (mecharv):

Yep now they seem right

OpenStudy (queenbb):

lol sorry

OpenStudy (daniyil_the_spy):

u will have to do substitution

OpenStudy (queenbb):

i dont know what you mren

OpenStudy (mecharv):

x+y=-1 3x-y=21 Multiply the first equation by 3 3x+3y=-3 3x-y=21 Now subtract both the equation, 4y=-24 y=-6 now which is the only option that has -6 for y?

OpenStudy (queenbb):

so it it (5,-6)

OpenStudy (daniyil_the_spy):

u will have to substitute Y with X to find x

OpenStudy (daniyil_the_spy):

I agree @mecharv

OpenStudy (queenbb):

so is (5-6) the answer

OpenStudy (daniyil_the_spy):

yes

OpenStudy (mecharv):

@queenbb Cheers :)